Consider a June Euro currency call with K = 1.025 and a call premium of 0.045. Suppose the current spot exchange rate of 1 Euro (Euro-USD) is 1.125 (1€ = $1.125). The breakeven Euro exchange rate is ________________. With the current € exchange rate, the intrinsic value of this Euro call is ______________, the time value is _____________, and the profit of the currency call is ________________.
Answer #1

Solution:

It is given that the strike price = 1.025

Call premium = 0.045

Spot price = 1.125

Part A )

Break-even point = strike price + premium paid = 1.025+0.045 = 1.07

Part B )

Intrinsic value = Spot price - strike price = 1.125-1.025 = 0.10

Part C)

Time value of option = option premium - Intrinsic value = 0.045 -0.1 =- 0.055

Time value is negative according to equation but it can't be less than zero hence time value = 0

Part D )

Profit = current spot price - break-even point = 1.125 - 1.070 = 0.018 = 0.055